# Accurate Screening of Functional Materials with Machine-Learning Potential and Transfer-Learned Regressions: Heusler Alloy Benchmark

### v1

@@ Abstract
A machine learning-accelerated high-throughput (HTP) workflow for the discovery of magnetic materials is presented. As a test case, we screened quaternary and all-$d$ Heusler compounds for stable compounds with large magnetocrystalline anisotropy energy ($E_{\mathrm{aniso}}$). Structure optimization and evaluation of formation energy and distance to hull convex were performed using the eSEN-30M-OAM interatomic potential, while local magnetic moments, phonon stability, magnetic stability, and $E_{\mathrm{aniso}}$ were predicted by eSEM models trained on our DxMag Heusler database. A frozen transfer learning strategy was employed to improve accuracy. Candidate compounds identified by the ML-HTP workflow were validated with density functional theory, confirming high predictive precision. We also benchmark the performance of different MLIPs, and discuss the fidelity of local magnetic moment prediction and its extension to other magnetic materials.
@@ 1 Introduction
The high-throughput (HTP) screening approach has emerged as a powerful strategy for accelerating the discovery of novel materials by systematically exploring large chemical spaces computationally or experimentally [1, 2, 3]. The density functional theory (DFT) based HTP workflows have been widely employed to identify materials with target properties. However, as the search space increases, the associated computational cost becomes unsustainable, often restricting screening efforts to a manageable subspace [4, 5]. To address this bottleneck, machine learning (ML) offers a promising route by drastically reducing computational costs. In this work, we demonstrate the robust integration of state-of-the-art ML techniques into the HTP workflow (ML-HTP) through a practical case study focused on screening quaternary and all-$d$ Heusler compounds for stable candidates with strong magnetic anisotropy energy ($E_{\mathrm{aniso}}$).
@@
Initial realizations of the ML-HTP paradigm relied on ML models that utilize compositional descriptors as input embeddings [6, 7, 8]. These models directly map chemical formulas to target properties, offering efficiency and simplicity. However, composition-based models are inherently unable to distinguish compounds with identical stoichiometry but different atomic arrangements. One workaround involves assigning layer indices to atomic sites, but this approach fixes the number of sites and can yield inconsistent predictions for symmetry-equivalent structures [9, 10]. Crystal graph based models do not have such drawbacks since they explicitly incorporate structure information as input, capturing structure-property relationships more accurately [11, 12]. However, crystal graph based models introduce an additional computational step, as geometry optimization must precede property prediction.
@@
Although a single DFT optimization typically requires only a few minutes, the cumulative cost of screening a large number of candidate compounds becomes prohibitively expensive, particularly for magnetic systems where multiple magnetic configurations must be considered. A promising solution lies in leveraging machine learning interatomic potentials (MLIPs), which can accelerate structure optimization by several orders of magnitude relative to DFT. The MLIP field has witnessed rapid advancements in recent years, with many crystal graph based models proposed. Despite the conceptual appeal, reliable and robust MLIP-based structure optimization has only become practical recently, especially following the release of the large-scale and diverse Meta Open Materials 2024 Dataset (OMat24) training dataset in 2024 [13]. This is demonstrated in the current work by benchmarking several MLIPs, ranging from early-stage implementations to state-of-the-art developments.
@@
With optimized structures, property can be predicted using machine learning regressor model (MLRM). This approach substantially reduces computational cost, particularly for properties that are expensive to compute via DFT, such as phonon spectra, conductivity, magnetic critical temperature ($T_{{\rm c}}$), and $E_{\mathrm{aniso}}$. However, training accurate MLRMs typically requires large, high-quality datasets. To overcome this challenge, transfer learning (TL) techniques can be employed to adapt pretrained ML models to new tasks [14, 15, 16]. TL leverages models that have already learned generalizable representations from extensive datasets and fine-tunes them using smaller, task-specific datasets. This strategy enhances predictive accuracy while substantially reducing data requirements.
@@
As a case study, we conducted a ML-HTP screening focused on Heusler compounds, which have garnered significant attention due to their diverse functional properties, technological potential, and the structure complexity [17]. Numerous DFT-HTP screenings have been carried out to identify candidates with different desirable properties [1, 4, 18, 5, 19, 20]. In our previous work, we developed DXMag Computational Heusler Database (HeuslerDB), a comprehensive database encompassing nearly all conventional ternary Heusler compounds. The present study significantly extended the search space to include quaternary and all-$d$ Heusler compounds, targeting stability and $E_{\mathrm{aniso}}$ as key screening criteria. In earlier DFT-HTP studies, 10 large $E_{\mathrm{aniso}}$ candidates were identified out of 286 selected compositions, and 15 among 29,784 Co-based structures [4, 5]. This low yield underscores the rarity of such materials and highlights the difficulty of this search problem, making it a stringent test case for ML-HTP approaches.
@@
In this work, we demonstrate the use of MLIP and TL-MLRM s as drop-in replacements for DFT structure optimization and property evaluation within HTP framework, as illustrated in Fig. 1(a). As a practical application, we employed this approach to identify conventional quaternary and all-$d$ Heusler compounds with large $E_{\mathrm{aniso}}$, while simultaneously satisfying thermodynamic, dynamic, and magnetic stability. Structure optimization and thermodynamic stability evaluation were performed using the eSEN-30M-OAM MLIP. The following predictions of local magnetic moment ($\{\boldsymbol{m}_{i}\}$), minimum phonon frequency ($\omega_{\mathrm{min}}$), $T_{{\rm c}}$, and $E_{\mathrm{aniso}}$ were performed using MLRM s. The MLRM s were trained via frozen transfer learning, using eSEN-30M-OAM MLIP as the base model and fine-tuned using HeuslerDB data and newly computed data. Identified candidates were validated through DFT calculations to demonstrate the significant reliability of this ML-HTP approach. We further examine key factors that influence the performance of such ML-HTP workflows, including accuracy of MLIP-based structure optimization, the magnetic configuration prediction, and the frozen transfer learning technique.
Figure 1: (a) Workflow overview of MLIP and TL-MLRM as a drop-in replacement for DFT in HTP screening. The TL-MLRM is initialized from the eSEN-30M-OAM MLIP, with the embedding layers and the first seven layers held fixed during training. The eSEN architecture schematic is adapted from the original publication introducing the model [21]. (b) Compositional space covered in the HTP screening. (c) Detailed HTP workflow for identifying conventional quaternary and all-$d$ Heusler compounds with stability and strong $E_{\mathrm{aniso}}$.
@@ 2 Results
To accumulate data for training $E_{\mathrm{aniso}}$ MLRM, we first computed the $E_{\mathrm{aniso}}$ of conventional ternary Heusler compounds within HeuslerDB using DFT. The $E_{\mathrm{aniso}}$ of some Heusler compounds were reported in the previous work and the agreement between our DFT results and previous work is demonstrated in the Supplementary Materials [4, 5]. Among all conventional ternary Heusler compounds, 2190 (7.9 %) exhibit $E_{\mathrm{aniso}}$ magnitude greater than 1 MJ/m3. When further screened for thermodynamic, dynamical, and magnetic stability, only 135 compounds (0.5 %) meet both the high $E_{\mathrm{aniso}}$ and stability criteria. These low percentages underscore the difficulty of identifying stable, high $E_{\mathrm{aniso}}$ compounds and highlight the need for more efficient screening methods demonstrated in current work. The compositional space covered in ML-HTP is shown in Fig. 1(b) and the ML-HTP workflow containing selection criteria is shown in Fig. 1(c). The detailed workflow and computation methods are described in Sec. 4.
@@ 2.1 Validation of ML-HTP selected candidates
Figure 2: Validation summary of ML-HTP approach. For candidate lists of conventional quaternary and all-$d$ Heusler compounds, the percentages that satisfy the screening criteria by DFT validation are shown (ML-HTP quat. and ML-HTP all-$d$). The precision of the ML models, as measured on the test set of conventional ternary compounds, is also reported (ML-HTP tern. test). To demonstrate efficiency improvement, the percentages of compounds in the whole conventional ternary Heusler space that meet the same criteria using DFT-HTP screening are included (DFT-HTP tern.).
@@
Using eSEN-30M-OAM MLIP in combination with MLRM s, we screened almost all conventional quaternary and all-$d$ ternary and quaternary Heusler spaces for stable compounds with high $E_{\mathrm{aniso}}$. Among 131,544 conventional quaternary Heusler and 104,139 all-$d$ Heusler compounds, the workflow yielded 366 and 924 candidates, respectively. To evaluate the reliability of this ML workflow, all candidates were validated using DFT calculations and the results are provided in the Supplementary Materials. The validation results are summarized in Fig. 2 as ML-HTP quat. and ML-HTP all-$d$ bins and the performances across $c/a$ ratio, formation energy ($\Delta E$), distance to convex hull ($\Delta H$), $\{\boldsymbol{m}_{i}\}$, $\omega_{\mathrm{min}}$, $T_{{\rm c}}$, and $E_{\mathrm{aniso}}$ are shown. For each property, the percentage of candidates meeting the screening criterion in DFT validation is reported. The precisions of the ML models, measured on the test set of conventional ternary compounds, are also provided as ML-HTP tern. test bins for comparison. To demonstrate the efficiency improvement, the percentages of compounds in the all conventional ternary Heusler space that satisfy the same criterion are included and marked as DFT-HTP tern..
@@
In ML-HTP, the $c/a$ ratio, $\Delta E$, and $\Delta H$ are obtained from optimized structure and the corresponding energy by eSEN-30M-OAM MLIP. Since a non-zero $E_{\mathrm{aniso}}$ requires the Heusler compound to adopt a tetragonal phase, we applied a screening threshold of |c/a1|>0.01|c/a-1|>0.01 to identify tetragonality. Notably, all ML-selected candidates remained tetragonal in DFT validation, confirming that eSEN-30M-OAM reliably distinguishes between cubic and tetragonal phases. A more detailed discussion of the prediction performance for lattice parameters $a$, $c$, and the $c/a$ ratio, and comparisons with other MLIP models, are provided in Sec. 2.3.
@@
Using energies of candidate compounds, elements, and competing phases predicted by eSEN-30M-OAM, the $\Delta E$ and $\Delta H$ were calculated. The criteria of $\Delta E$ <0<0 eV/atom and $\Delta H$ <0.22<0.22 eV/atom were employed to identify thermodynamically stable candidates, following the thresholds established in our previous DFT-HTP study [20]. Among ML-selected candidates, 99.1% of conventional quaternary and 97.8% of all-$d$ Heusler compounds were validated to have $\Delta E_{\mathrm{DFT}}<0$ eV/atom. Similarly, 96.4% (quaternary) and 98.8% (all-$d$) of the compounds were found to have $\Delta H_{\mathrm{DFT}}<0.22$ eV/atom. These high validation rates demonstrate that state-of-the-art MLIPs, such as eSEN-30M-OAM, can reliably assess thermodynamic stability.
@@
It is important to note that the eSEN-30M-OAM MLIP used is a general-purpose, pretrained model without any fine-tuning specific to the Heusler chemical space. Thus, these results highlight its strong generalization, making it an effective drop-in replacement for DFT-based optimization and thermodynamic stability assessment in HTP workflow. Notably, the model achieves strong performance on the studied magnetic systems, despite not explicitly incorporating magnetic moments into its architecture or training. This strong performance and generalization is particularly valuable for the initial screening of novel material systems, where MLIPs can greatly reduce the search space by rapidly and reliably estimating optimized structures and thermodynamic stability.
@@
The properties $\{\boldsymbol{m}_{i}\}$, $\omega_{\mathrm{min}}$, $T_{{\rm c}}$, and $E_{\mathrm{aniso}}$ were predicted using MLRM s applied to MLIP-optimized structures. Because $E_{\mathrm{aniso}}$ is a magnetic property, $\{\boldsymbol{m}_{i}\}$ was first predicted, and a screening threshold of $\sum|\boldsymbol{m}_{i}|>0.1$ $\mu_{\mathrm{B}}$/f.u was applied to identify magnetic compounds. DFT validations confirmed that this classification achieved 100% precision. Moreover, the $\{\boldsymbol{m}_{i}\}$ MLRM accurately predicts both the magnitude and sign of local moments, as discussed in detail in Sec. 2.5. Magnetic system identification is a critical yet computationally demanding step in DFT-HTP, as multiple initial $\{\boldsymbol{m}_{i}\}$ values must be tested and the low fraction of magnetic systems in some material families can lead to substantial wasted computation. By incorporating the $\{\boldsymbol{m}_{i}\}$ MLRM method as a pre-screening step, the search space can be substantially reduced.
@@
Relaxing the screening thresholds increases the pool of candidate compounds and captures promising cases missed initially, though at the expense of reduced precision. Nonetheless, the curated list of compounds that meet stability criteria, and further magnetic system criterion, serves as an efficient starting point for investigation of additional functional properties. For readers interested in exploring an expanded candidate list, the full set of ML-HTP predicted data of 113,544 conventional quaternary and 104,139 all-$d$ Heusler compounds will be added to the HeuslerDB, which is publicly available at https://www.nims.go.jp/group/spintheory/.

### v2

@@ Abstract
We present a machine learning-accelerated high-throughput (HTP) workflow for the discovery of functional materials. As a test case, quaternary and all-$d$ Heusler compounds were screened for stable compounds with large magnetocrystalline anisotropy energy ($E_{\mathrm{aniso}}$). Structure optimization and evaluation of formation energy and energy above the convex hull were performed using the eSEN-30M-OAM interatomic potential, while local magnetic moments, phonon stability, magnetic stability, and $E_{\mathrm{aniso}}$ were predicted by eSEN models trained on our DxMag Heusler database. A frozen transfer learning strategy was employed to improve accuracy. Candidate compounds identified by the ML-HTP workflow were validated with density functional theory, confirming high predictive precision. We also benchmark the performance of different uMLIPs, discuss the fidelity of local magnetic moment prediction, and demonstrate generalization to unseen elements via transfer learning from a universal interatomic potential.
@@ 1 Introduction
The high-throughput (HTP) screening approach has emerged as a powerful strategy for accelerating the discovery of novel materials by systematically exploring large chemical spaces computationally or experimentally [sanvitoAcceleratedDiscoveryNew2017, zhangHighthroughputDesignMagnetic2021, barwalLargeMagnetoresistanceHigh2024]. The density functional theory (DFT)-based HTP workflows have been widely employed to identify materials with target properties. However, as the search space increases, the associated computational cost becomes unsustainable, often restricting screening efforts to a manageable subspace [faleevHeuslerCompoundsPerpendicular2017, huHighthroughputDesignCobased2023]. To address this bottleneck, machine learning (ML) offers a promising route by drastically reducing computational costs. In this work, we demonstrate the robust integration of state-of-the-art ML techniques into the HTP workflow (ML-HTP) through a practical case study focused on screening quaternary and all-$d$ Heusler compounds for stable candidates with strong magnetic anisotropy energy ($E_{\mathrm{aniso}}$).
@@
Initial realizations of the ML-HTP paradigm relied on ML models that utilize compositional descriptors as input features [hilgersMachineLearningbasedEstimation2025, baigutlinMachineLearningAlgorithms2024, mitraMachineLearningApproach2023, liuMachineLearningPredict2021]. These models directly map chemical formulas to target properties, offering efficiency and simplicity. However, composition-based models are inherently unable to distinguish compounds with identical stoichiometry but different atomic arrangements. One workaround involves assigning layer indices to atomic sites, but this approach fixes the number of sites and can yield inconsistent predictions for symmetry-equivalent structures [xieScreeningNewQuaternary2023, luExplainableAttentionCNN2025]. Crystal graph-based models do not have such drawbacks since they explicitly incorporate structure information as input, capturing structure-property relationships more accurately [xieCrystalGraphConvolutional2018a, choudharyAtomisticLineGraph2021]. However, crystal graph-based models introduce an additional computational step, as geometry optimization must precede property prediction.
@@
Although a single DFT optimization typically requires only a few minutes, the cumulative cost of screening a large number of candidate compounds becomes prohibitively expensive, particularly for magnetic systems where multiple magnetic configurations must be considered. A promising solution lies in leveraging universal machine learning interatomic potentials (uMLIPs), which can accelerate structure optimization by several orders of magnitude relative to DFT. The uMLIP field has witnessed rapid advancements in recent years, with many crystal graph-based models proposed. Despite the conceptual appeal, reliable and robust uMLIP-based structure optimization has only become practical recently, especially following the release of the large-scale and diverse Meta Open Materials 2024 Dataset (OMat24) training dataset in 2024 [barroso-luqueOpenMaterials20242024]. This is demonstrated in the current work by benchmarking several uMLIPs, ranging from early-stage implementations to state-of-the-art developments.
@@
With optimized structures, properties can be predicted using machine learning regression model (MLRM). This approach substantially reduces computational cost, particularly for properties that are expensive to compute via DFT, such as phonon spectra, conductivity, magnetic critical temperature ($T_{{\rm c}}$), and $E_{\mathrm{aniso}}$. However, training accurate MLRMs typically requires large, high-quality datasets. To overcome this challenge, transfer learning (TL) techniques can be employed to adapt pretrained ML models to new tasks [yamadaPredictingMaterialsProperties2019, leeTransferLearningMaterials2021a, hoffmannTransferLearningLarge2023]. TL leverages models that have already learned generalizable representations from extensive datasets and fine-tunes them using smaller, task-specific datasets. This strategy enhances predictive accuracy while substantially reducing data requirements.
@@
As a case study, we conducted a ML-HTP screening on Heusler compounds, which have garnered significant attention due to their diverse functional properties, technological potential, and structure complexity [heComputationallyAcceleratedDiscovery2022]. Numerous DFT-HTP screenings have been carried out to identify candidates with different desirable properties [sanvitoAcceleratedDiscoveryNew2017, faleevHeuslerCompoundsPerpendicular2017, nokyGiantAnomalousHall2020, huHighthroughputDesignCobased2023, xingChemicalsubstitutiondrivenGiantAnomalous2024a, xiaoHighthroughputComputationalScreening2025]. In our previous work, we developed DXMag Computational Heusler Database (HeuslerDB), a comprehensive database encompassing nearly all conventional ternary Heusler compounds. The present study significantly extended the search space to include quaternary and all-$d$ Heusler compounds, targeting stability and $E_{\mathrm{aniso}}$ as key screening criteria. In earlier DFT-HTP studies of Heusler compounds, 10 candidates with large $E_{\mathrm{aniso}}$ were identified out of 286 selected compositions, and 15 among 29,784 Co-based structures [faleevHeuslerCompoundsPerpendicular2017, huHighthroughputDesignCobased2023]. This low yield underscores the rarity of such materials and highlights the difficulty of this search problem, making it a stringent test case for ML-HTP approaches.
Previous studies attempted ML approaches to $E_{\mathrm{aniso}}$ in systems outside the Heusler family. They employed early ML models, such as crystal graph convolutional neural networks (CGCNN) and compositional-descriptor models, to predict $E_{\mathrm{aniso}}$ in Fe–Co–N alloys and physics-informed 2D materials, working on the order of hundreds of compounds [xieDataDrivenStudiesMagnetic2021, liaoPredictingMagneticAnisotropy2022, duttaMachineLearningAssisted2022]. Here, we employ state-of-the-art ML methods to extend the scope to hundreds of thousands of compounds with improved accuracy and practicality.
@@
In this work, we demonstrate the use of uMLIP and TL-MLRM s as drop-in replacements for DFT structure optimization and property evaluation within HTP framework, as illustrated in Fig. 1(a). As a practical application, we employed this approach to identify conventional quaternary and all-$d$ Heusler compounds with large $E_{\mathrm{aniso}}$, while simultaneously satisfying thermodynamic, dynamic, and magnetic stability. Structure optimization and thermodynamic stability evaluation were performed using the eSEN-30M-OAM uMLIP [fuLearningSmoothExpressive2025]. The following predictions of local magnetic moment ($\{\boldsymbol{m}_{i}\}$), minimum phonon frequency ($\omega_{\mathrm{min}}$), $T_{{\rm c}}$, and $E_{\mathrm{aniso}}$ were performed using MLRM s. The MLRM s were trained via frozen transfer learning, using eSEN-30M-OAM uMLIP as the base model and fine-tuned using HeuslerDB data and newly computed data. ML-selected candidates were validated through DFT calculations to demonstrate the significant reliability of this ML-HTP approach. We further examine key factors that influence the performance of such ML-HTP workflows, including accuracy of uMLIP-based structure optimization, the magnetic configuration prediction, the performance of frozen transfer learning technique and generalization to unseen elements. The underlying code for the ML-HTP workflow is made available as open-source packages MLIP High-throughput Optimization and Thermodynamics (MLIP-HOT) and MLIP Frozen Transfer Learning (MLIP-FTL), which can be found on our group’s website and Git repository.
Figure 1: Frozen transfer learning overview and ML-HTP workflow. a) Schematic of the development of the MLRM via frozen transfer learning, using eSEN-30M-OAM uMLIP as the base model. The uMLIP is used to perform structure optimization, formation energy calculation, and convex hull distance evaluation. The MLRM predicts properties from structures. b) Workflow of the case study which identified stable conventional quaternary and all-$d$ Heusler compounds exhibiting strong $E_{\mathrm{aniso}}$. Counts of quaternary and all-$d$ compounds at each stage are reported as quaternary/all-$d$.
@@ 2 Results
To accumulate data for training $E_{\mathrm{aniso}}$ MLRM, we first computed the $E_{\mathrm{aniso}}$ of conventional ternary Heusler compounds within HeuslerDB using DFT. The $E_{\mathrm{aniso}}$ of some Heusler compounds were reported in the previous work and the agreement between our DFT results and previous work is demonstrated in Fig. S1 [faleevHeuslerCompoundsPerpendicular2017, huHighthroughputDesignCobased2023]. Among all conventional ternary Heusler compounds, 2190 (7.9 %) exhibit an $E_{\mathrm{aniso}}$ magnitude greater than 1 MJ/m3. When further screened for thermodynamic, dynamical, and magnetic stability, only 135 compounds (0.5 %) meet both the high $E_{\mathrm{aniso}}$ and stability criteria, which are presentated in Table S2. These low percentages underscore the difficulty of identifying stable, high $E_{\mathrm{aniso}}$ compounds and highlight the need for more efficient screening methods demonstrated in current work as a case study. The ML-HTP workflow for this case study is summarized in Fig. 1(b); detailed computational procedures are provided in Sec. 4.
For conventional quaternary compounds, we enumerated all combinations where $X$ and $Y$ are transition metals from the $d$-block (excluding Tc and Hg), and $Z$ is a main-group element from groups 13, 14, or 15 of the $p$-block. In addition, La and Lu were included for $X$ and $Y$ because their 4ff orbitals are either empty or fully filled. This exhaustive enumeration, accounting for symmetry constraints, yielded 131,544 unique compositions. For the all-$d$ Heuslers, we extended the screening space to include $d$-block transition metals together with La and Lu across all four sites ($X_{1}$, $X_{2}$, $Y$, and $Z$), resulting in a separate set of 105,763 unique compositions. A schematic of the screened chemical space is presented in Supplementary Information as Fig. S2.
@@ 2.1 Validation of ML-HTP selected candidates
Figure 2: DFT validation summary of ML-HTP selected compounds. For ML-selected candidate lists of conventional quaternary (334) and all-$d$ (924) Heusler compounds, the percentages that DFT results satisfy the screening criteria (i.e., the ML‑HTP precision) are shown as blue and yellow bars. For comparison, the precision of the ML models measured on the test set of conventional ternary compounds is also shown as green bars. The test set size for $c/a$ ratio, $\Delta E$, and $\Delta H$ is 10,000 and for $\{\boldsymbol{m}_{i}\}$, $\omega_{\mathrm{min}}$, $T_{{\rm c}}$, and $E_{\mathrm{aniso}}$ is 10% of the dataset size shown in Table 1.
@@
Using eSEN-30M-OAM uMLIP in combination with MLRM s, we screened almost all conventional quaternary and all-$d$ Heusler spaces for stable compounds with high $E_{\mathrm{aniso}}$. As a result, 334 and 924 candidates were found, respectively. To evaluate the reliability of this ML workflow, all candidates were validated using DFT calculations. The results are summarized in Fig. 2 and detailed data are provided in Tables S3 and S4. The percentages that DFT results satisfy the screening criteria (i.e., precision) are shown as blue and yellow bars for conventional quaternary and all-$d$ Heusler compounds, respectively. The selection criteria include $c/a$ ratio, $\Delta E$, $\Delta H$, $\{\boldsymbol{m}_{i}\}$, $\omega_{\mathrm{min}}$, $T_{{\rm c}}$, and $E_{\mathrm{aniso}}$. For comparison, the precisions of the ML models, measured on the test set of conventional ternary compounds, are also provided as green bars.
@@
In ML-HTP, the $c/a$ ratio, $\Delta E$, and $\Delta H$ are obtained from the optimized structure and the corresponding energy by eSEN-30M-OAM uMLIP. During the structure optimization process, relaxations were performed starting from multiple initial structures, and the relaxed structure with the lowest energy was selected. Since a non-zero $E_{\mathrm{aniso}}$ requires the Heusler compound to adopt a tetragonal phase, we applied a screening threshold of |c/a1|>0.01|c/a-1|>0.01 to identify tetragonality. Notably, all ML-selected candidates remained tetragonal in DFT validation, confirming that eSEN-30M-OAM reliably distinguishes between cubic and tetragonal phases. A more detailed discussion of structure optimization and performance for lattice parameters $a$, $c$, and the $c/a$ ratio prediction, along with performance of other uMLIP s, are provided in Sec. 2.3.
@@
Using energies of candidate compounds, elements, and competing phases predicted by eSEN-30M-OAM, the $\Delta E$ and $\Delta H$ were calculated. The criteria of $\Delta E$ <0<0 eV/atom and $\Delta H$ <0.22<0.22 eV/atom were employed to identify thermodynamically stable candidates, following the thresholds established in our previous DFT-HTP study [xiaoHighthroughputComputationalScreening2025]. Among ML-selected candidates, 99.1% of conventional quaternary and 97.8% of all-$d$ Heusler compounds were validated to have $\Delta E_{\mathrm{DFT}}<0$ eV/atom. Similarly, 96.4% (quaternary) and 98.8% (all-$d$) of the compounds were found to have $\Delta H_{\mathrm{DFT}}<0.22$ eV/atom. These high validation rates demonstrate that state-of-the-art uMLIPs, such as eSEN-30M-OAM, can reliably assess thermodynamic stability.
@@
It is important to note that the eSEN-30M-OAM uMLIP used is a general-purpose, pretrained model without any fine-tuning specific to the Heusler chemical space. Thus, these results highlight its strong generalization, making it an effective drop-in replacement for DFT-based optimization and thermodynamic stability assessment in HTP workflow. Notably, the model achieves strong performance on the studied magnetic systems, despite not explicitly incorporating magnetic moments into its architecture or training. This strong performance and generalization are particularly valuable for the initial screening of novel material systems, where uMLIPs can greatly reduce the search space by rapidly and reliably estimating optimized structures and thermodynamic stability.
@@
The properties $\{\boldsymbol{m}_{i}\}$, $\omega_{\mathrm{min}}$, $T_{{\rm c}}$, and $E_{\mathrm{aniso}}$ were predicted using MLRM s applied to uMLIP-optimized structures. Because $E_{\mathrm{aniso}}$ is a magnetic property, $\{\boldsymbol{m}_{i}\}$ was predicted, and a screening threshold of $\sum|\boldsymbol{m}_{i}|>0.1$ $\mu_{\mathrm{B}}$/f.u was applied to identify magnetic compounds. DFT validations confirmed all ML-selected candidates to be magnetic. Moreover, the $\{\boldsymbol{m}_{i}\}$ MLRM accurately predicts both the magnitude and sign of local moments, as discussed in detail in Sec. 2.5. Magnetic system identification is a critical yet computationally demanding step in DFT-HTP, as multiple initial $\{\boldsymbol{m}_{i}\}$ values must be tested and the low fraction of magnetic systems in some material families can lead to substantial wasted computation. By incorporating the $\{\boldsymbol{m}_{i}\}$ MLRM method as a pre-screening step, the search space can be substantially reduced.
@@
Relaxing the screening thresholds increases the pool of candidate compounds and might capture promising cases missed initially at the expense of more false positives. Additionally, the curated list of compounds that meet stability criteria, and further magnetic system criterion, serves as an efficient starting point for investigation of other functional properties. For readers interested in exploring an expanded candidate list, the full set of ML-predicted data for 131,544 conventional quaternary and 105,763 all-$d$ Heusler compounds will be accessible through HeuslerDB.